OpenStudy (anonymous):

What is the circumference of the circle whose equation is (x-9)^2+(y-3)^2=64?

OpenStudy (anonymous):

a) 16pi b)4pi c)8pi d)64pi

OpenStudy (zzr0ck3r):

radius is 8 circumference is 2*r*pi

OpenStudy (zzr0ck3r):

what is the answer?

OpenStudy (anonymous):

See the basic equation of a circle with radius 'r' is x^2 + y^2 = r^2 And this is a modification of it (x-a)^2 + (y-b)^2 = r^2 but the radius of circle 'r' is still the same. In a way the circle is just moved, with (a,b) being new centre points(Previously they were 0,0). Anyway, the circumference of a circle = 2pi(r) There for here (x-9)^2+(y-3)^2=64= 8^2, r=8 So circumference = 2pi(8)= 16pi Got it?

OpenStudy (zzr0ck3r):

I e what is 2*8

OpenStudy (anonymous):

so 16pi is right?

OpenStudy (anonymous):

yeah, it is.

OpenStudy (anonymous):

thanks all

OpenStudy (anonymous):

you got what i said?

OpenStudy (anonymous):

yes =)

OpenStudy (anonymous):

Good. :)